OpenStudy (anonymous):

A grid shows the positions of a subway stop and your house.The subway stop is located at (7, -6) and your house is located at (-9,5). What is the distance to the nearest unit between your house and the subway stop?

OpenStudy (anonymous):

(9-2)^2+(-9-(-5))^2=c^2 65=c^2 closest unit is 8